A test is given to a class of 3 students. Marks obtained are 2,4 and 6, assume the students constitute the population. Find the standard deviation of the population

List all possible samples of size 2 that can drawn with replacement from the population

Determine the mean of each sample

Determine the mean of the sample means

Determine the standard deviation of sample means


1
Expert's answer
2021-12-06T16:12:44-0500

Lets list all such samples and calculate their means

(2, 4) "m={\frac {2+4} 2}=3"

(4, 2) "m={\frac {2+4} 2}=3"

(2, 6) "m={\frac {2+6} 2}=4"

(6, 2) "m={\frac {2+6} 2}=4"

(4, 6) "m={\frac {6+4} 2}=5"

(6, 4) "m={\frac {6+4} 2}=5"


Lets calculate M - mean of the sample means

"M={\frac{2*3+2*4+2*5} 6}=4"


Lets calculate "\sigma" - the standard deviation of the sample means

"\sigma=\sqrt{{\frac {2(3-4)^2+2(4-4)^2+2*(5-4)^2} {6-1}}}=\sqrt{\frac 4 5}\approx 0.89"